### v1

@@ Abstract
Since its release in November 2022, ChatGPT has shaken up Stack Overflow, the premier platform for developers’ queries on programming and software development. Demonstrating an ability to generate instant, human-like responses to technical questions, ChatGPT has ignited debates within the developer community about the evolving role of human-driven platforms in the age of generative AI. Two months after ChatGPT’s release, Meta released its answer with its own Large Language Model (LLM) called LLaMA: the race was on. We conducted an empirical study analyzing questions from Stack Overflow and using these LLMs to address them. This way, we aim to ① measure user engagement evolution with Stack Overflow over time; ② quantify the reliability of LLMs’ answers and their potential to replace Stack Overflow in the long term; ③ identify and understand why LLMs fails; and ④ compare LLMs together. Our empirical results are unequivocal: ChatGPT and LLaMA challenge human expertise, yet do not outperform it for some domains, while a significant decline in user posting activity has been observed. Furthermore, we also discuss the impact of our findings regarding the usage and development of new LLMs.
@@
ChatGPT LLaMa Stack Overflow Empirical study Reliability
journal: Empirical Software Engineering
@@ 1 Introduction
During software development, practitioners are expected to adopt different methods and approaches to support their work. These involve making technical and non-technical decisions, encompassing the selection of programming languages, frameworks, and methodologies, and the prominent impact of the taken decisions on the software development process (Yli-Huumo et al., 2016; Tamburri et al., 2013; Dias et al., 2020). When it comes to technical support, practitioners have become familiar with leveraging online Question and Answer (Q&A) web forums as instrumental aids, playing an essential role in different aspects, like API learning, code compression, and fixing or related problem information  (Rubei et al., 2020; Squire, 2015). Notably, Stack Overflow  [^1] https://stackoverflow.com/ stands out as a preeminent online hub within the technology community (Xia et al., 2017). However, owing to the inherent nature of these Q&A platforms, which are contingent upon the accumulation of both individual and collective human knowledge, the potential for errors and the propagation of inaccurate information remains a concern (Verdi et al., 2020; Ragkhitwetsagul et al., 2019; Zhang et al., 2018).
@@
With the advance of Large Language Models (LLMs), practitioners and researchers have explored their potential to perform different tasks, like pair programming assistant, simulating human behavior (questionnaire respondents), and even empirical study supporter  (Dakhel et al., 2023; Hämäläinen et al., 2023; Liang et al., 2023; Hou et al., 2023). LLMs exhibit a capacity to prompt for a variety of subjects and instantaneously provide comprehensive and concise explanations (Touvron et al., 2023). Furthermore, with the release of ChatGPT  [^2] https://openai.com/chatgpt, publicly available for the general audience, users have prompted them for different purposes, like simulating a Q&A online web forum. As a result, these models, by engaging in human-like tasks, can continuously learn from user feedback and be further explored by fine-tuning and improving accuracy (Bakker et al., 2022). Recently, researchers observed a decline in the number of answers to posted questions related to different programming languages, even before the release of LLMs (Blanco et al., 2020; Syam et al., 2023). Although Stack Overflow adopts different approaches to keep users engaged, like offering badges (Lu et al., 2021), the adoption of LLMs might impact user engagement as well.
@@
Therefore, practitioners might turn to using more and recurrently LLMs instead of checking truthful and verifiable sources. Nonetheless, exclusively relying on information generated by LLMs is a threat, as LLMs might generate inaccurate information (Goodrich et al., 2019). To overcome the imminent risks, Stack Overflow has proactively announced prohibiting the use of information generated by LLMs (StackOverflow, 2023a). However, considering the difficulty of detecting contents generated by LLMs (Tang et al., 2023; Sadasivan et al., 2023), they later announced OverflowAI, an integration of generative AI into the platform impacting users, products, and IDEs (StackOverflow, 2023b). Even under these new circumstances, there remains the need to evaluate the reliability of the generated answers and to what extent LLMs can be leveraged.
@@
Such concern about the reliability of LLMs is becoming a general and recurrent topic, also targeted by other research areas. For example, Lee et al. (2023) investigate the reliability of contents generated by ChatGPT in medicine. Based on scenario examples of potential medical use, the authors evaluate the current abilities of ChatGPT under different situations. The model performs satisfactorily, though the authors also observe some hallucinations, like inferring information unrelated to previously shared information. The authors further explore these hallucinations in ChatGPT 4, showing that the model could catch them.
@@
In the same way, evaluating the reliability of LLMs in Software Engineering while assisting developers with their daily tasks is a relevant and important topic, bringing the attention of previous studies to this topic. Kabir et al. (2023) and Liu et al. (2023) investigate the correctness of ChatGPT by comparing the answers reported with human ones. However, both studies focus only on ChatGPT, while they do not deeply explore the most challenging domains faced by the model. Given LLMs’ current diversity and availability, it is important to investigate reliability and compare findings across different models. Furthermore, investigating the impact of LLMs on Stack Overflow user engagement and the different contexts in which these models are used is an open area not explored yet.
@@
To fill this knowledge gap, we explore LLMs’ impact, reliability, and challenges in this work, focusing on user engagement and the generation of answers for tech community practitioners. By reliability, we aim to investigate the power of LLMs to report trustful and valuable answers for the evaluated questions. To that end, we perform an empirical study analyzing questions and associated answers mined from Stack Overflow. We mine questions and related information from Stack Overflow for ten months (before and after ChatGPT’s release). Next, we select a representative number of the previously selected questions to prompt LLMs and then assess the reliability by comparing the generated answers with the original ones provided on Stack Overflow. Trying to mitigate previous studies’ diversity issues regarding LLMs under analysis, we investigate two models: ChatGPT-3.5 and LLaMA-2. While ChatGPT has been largely explored by practitioners and researchers, LLaMA represents an alternative option for similar proposes. Although both models are publicly available, LLaMA provides further support to users without additional costs (ChatGPT can be quite expensive) by allowing them to load and prompt the model locally.
@@
Our results show a continuous significant decline in Stack Overflow activity after ChatGPT’s release, which includes posting questions, answers, and comments. When prompting LLMs to address previous questions posted on Stack Overflow, we observe that ChatGPT-3.5 significantly outperforms LLaMA-2 related to textual similarity. Furthermore, we also observe a significant correlation between the number of tags associated with a question and the achieved textual similarity for ChatGPT. Although LLaMA does not present significantly better results than ChatGPT, this model represents an alternative, free, and public option for the tech community. Overall, the LLMs evaluated here perform well in most domains, with some challenges regarding questions targeting Frameworks and Libraries, leading to different observed impacts on specific topics. Although ChatGPT presents a more neutral sentiment when generating answers, LLaMA stands out with a more positive behavior.
@@
  • An empirical study evaluating the reliability of answers provided by ChatGPT and LLaMA in comparison to human-generated answers on Stack Overflow.
  • We measure and analyze user engagement trends with Stack Overflow over time and show that, since ChatGPT’s release, Stack Overflow shows a decline in use in general.
  • We identify and further explore the domains where LLMs fall short and show that for some topics, no statistical decline was observed in its associated questions.
@@
  • A dataset of questions from Stack Overflow and associated scripts available through our online AppendixOnline Appendix (2023):
@@ 2 Background
Stack Overflow: Stack Overflow (SO) is a platform where developers can ask questions and share their knowledge with others (Barua et al., 2014). It is designed to provide accurate answers to specific programming problems, fostering a community of learning and sharing among developers. Over time, developers have adopted this platform to discuss different topics, while building general knowledge with the tech community.
Large Language Model: A large language model (LLM) is a machine learning model trained on extensive sets of textual data to generate human-like text (Xu et al., 2022). With millions to billions of parameters, LLMs are used for various natural language processing tasks. Different LLMs have been proposed for tasks associated with software engineering, like StarCoderLi et al. (2023b), Copilot, and CodeBERT Feng et al. (2020).
ChatGPT: ChatGPT is a language model developed by OpenAI. It is designed to generate human-like text based on the prompts it receives. Built on the GPT architecture, it can be used for a variety of applications, among which are code generation, summarization, translation, testing, and documentation (Zheng et al., 2023; Ozkaya, 2023). Currently, OpenAI offers different models, but in this study, we focus on ChatGPT-3.5. The motivation for this decision is driven by the recurrent adoption of this version by related work; furthermore, it is the default version that users have access to through the app option without further costs.
LLaMA: LLaMA, released by Meta AI, is a series of large language models that come with different sizes (e.g., 7, 13, and 70 billion parameters for LLaMA-2) (Touvron et al., 2023). Similar to ChatGPT, it spans various use cases, such as interacting with humans. In our study, we relied on the latest LLaMA version, i.e., LLaMA-2. This version has similarities with ChatGPT, like its architecture based on transformers and its generative capabilities, leading to generate coherent and contextually relevant text based on input prompts. Although this model has not been explored like ChatGPT, we aim to investigate how these models differ and complement each other. Specifically for this study, we consider option LLaMA-2-7b as it represents the most accessible option for users, based on the required resources to load and use the model locally.
@@ 3 Study Setup
Our experimental methodology investigates the impact and reliability of LLMs on Stack Overflow. To this end, our study aims to answer the following research questions:
[…]
  • RQ2. How does the reliability of answers generated by ChatGPT and LLaMA-2 compare to those provided by Stack Overflow users? We aim to explore the power of LLMs to address questions previously answered by Stack Overflow users and compare the reliability of their generated answers.We randomly select a representative number of previously collected questions and prompt the models to answer them. Next, we compute the reliability by checking the textual and semantic similarity of the generated and original answers.
  • RQ3. What specific domains pose challenges for ChatGPT and LLaMA-2 when generating reliable and accurate answers? We investigate how to leverage LLMs’ adoption for practitioners. This way, based on the results of the previous RQ, we explore the factors that correlate with high or low similarity. For instance, based on the length of the question description LLMs can be fed with, we may assess whether generated answers have a higher or lower similarity. Next, we analyze the questions and extract their associated domains; in this study, we understand domain as the general topic associated with a given question. Based on the domains previously reported by Barua et al. (2014), we analyze the questions trying to classify them. Furthermore, we perform a sentiment analysis of the generated answers. Knowing that answers avoiding negative comments are more likely to be accepted, we aim to assess how close the generated answers would be to being accepted (Calefato et al., 2015).
@@
  • RQ4. How the release of ChatGPT change the profile of Stack Overflow users? To answer this RQ, initially, we compare the activeness of users on Stack Overflow pre- and post-release, considering posting questions, answers, and comments. Next, based on the different groups of users observed, we further explore how users arrange ChatGPT and Stack Overflow. Based on our findings, we discuss the challenges Stack Overflow faces regarding their new users while understanding their new current needs.
  • RQ5. How different domains from Stack Overflow are impacted by the release of ChatGPT? In this RQ, we investigate the impact of ChatGPT on specific domains of Stack Overflow. For that, we filter previously mined questions, which represent challenges for LLMs, based on our findings and compare the frequency of posted questions referring to these topics (tags). Based on our findings, we discuss possible factors that might influence the results and the different strategies for adopting/arranging Stack Overflow and LLMs from now on.
@@
The remaining of this section describes the empirical setup of our study. We structure our study in three main steps, as presented in Figure 1: ❶ we mine SO questions using its API from July 2022 till May 2023; then, we filter the questions with accepted answers and sort them based on their score. ❷ we select a representative number of questions from this subgroup and use them to prompt ChatGPT and LLaMA-2; and ❸ we further collect metrics regarding question contents and associated information.

### v2

@@ Abstract
Since its release in November 2022, ChatGPT has shaken up Stack Overflow, the premier platform for developers’ queries on programming and software development. Demonstrating an ability to generate instant, human-like responses to technical questions, ChatGPT has ignited debates within the developer community about the evolving role of human-driven platforms in the age of generative AI. Two months after ChatGPT’s release, Meta released its answer with its own Large Language Model (LLM) called LLaMA: the race was on. We conducted an empirical study analyzing questions from Stack Overflow and using these LLMs to address them. This way, we aim to ① quantify the reliability of LLMs’ answers and their potential to replace Stack Overflow in the long term; ② identify and understand why LLMs fail; ③ measure users’ activity evolution with Stack Overflow over time; and ④ compare LLMs together. Our empirical results are unequivocal: ChatGPT and LLaMA challenge human expertise, yet do not outperform it for some domains, while a significant decline in user posting activity has been observed. Furthermore, we also discuss the impact of our findings regarding the usage and development of new LLMs and provide guidelines for future challenges faced by users and researchers.
@@
ChatGPT , LLaMa , Stack Overflow , Empirical study , Reliability
journal: Journal of Systems and Software
@@ 1 Introduction
Practitioners must adopt different methods and approaches to support their work during software development. This involves making technical and non-technical decisions, encompassing the selection of programming languages, frameworks, and methodologies. These decisions have a prominent impact on the software development process [1, 2, 3]. When it comes to technical support, practitioners have become familiar with leveraging online Question and Answer (Q&A) web forums as instrumental aids playing an essential role in different aspects, like API learning, code compression, and fixing or getting related problem information  [4, 5]. Notably, Stack Overflow  [^1] https://stackoverflow.com/ stands out as a preeminent online hub within the technology community [6]. However, owing to the inherent nature of these Q&A platforms, which are contingent upon the accumulation of both individual and collective human knowledge, the potential for errors and the propagation of inaccurate information remains a concern [7, 8, 9].
@@
With the advance of Large Language Models (LLMs), practitioners and researchers have explored the potential of leveraging LLMs for different software engineering tasks; for example, pair programming assistants and simulating human behavior as questionnaire respondents  [10, 11, 12, 13]. Such a variety of usage is motivated by the capability of LLMs to provide comprehensive and concise explanations for diverse subjects instantaneously [14]. Furthermore, with the release of ChatGPT  [^2] https://openai.com/chatgpt, publicly available for the general audience, users have prompted LLMs for different purposes, like simulating a Q&A online web forum. As a result, these models, by engaging in human-like tasks, can continuously learn from user feedback (e.g., via fine-tuning) and improve their accuracy [15].
@@
Therefore, practitioners might turn to using more and recurrently LLMs instead of checking truthful and verifiable sources. Nonetheless, exclusively relying on information generated by LLMs is a threat, as LLMs might generate inaccurate information [16]. To overcome the imminent risks, Stack Overflow has proactively announced prohibiting the use of information generated by LLMs [17]. However, considering the difficulty of detecting contents generated by LLMs [18, 19], they later announced OverflowAI, an integration of generative AI into the platform impacting users, products, and IDEs [20]. Even under these new circumstances, it is still necessary to assess the generated answers and determine the extent to which LLMs can be effectively utilized.
@@
Such concern is becoming a general and recurrent topic, also targeted by other research areas. For example, Lee et al. [21] investigate the reliability of content generated by ChatGPT-3.5 in medicine. Based on scenario examples of potential medical use, the authors evaluate the current abilities of ChatGPT under different situations. The model performs satisfactorily, though the authors also observe some hallucinations, like inferring information unrelated to previously shared content. These hallucinations were further explored in ChatGPT 4 when the model could catch them.
@@
In the same way, evaluating the reliability of LLMs in Software Engineering while assisting developers with their daily tasks is a relevant and important topic, bringing the attention of previous studies. Kabir et al. [22] and Liu et al. [23] investigate the correctness of ChatGPT by comparing the answers reported with human ones. However, by reliability, we refer to the ability of LLMs to consistently produce correct and accurate behavior across time and varying conditions. While correctness is a key aspect of evaluating LLMs, reliability encompasses more than just factual accuracy. LLMs might provide correct answers, but they could also share content that, although technically correct, may still be misleading or harmful to users in some contexts. Therefore, evaluating such an aspect is crucial for ensuring LLMs do not produce misleading, harmful, or incomplete information, even when their responses are technically correct. Furthermore, previous studies focus only on ChatGPT, while they also do not deeply explore the most challenging domains faced by the model and how users deal with them. Given LLMs’ current diversity, different number of parameters supported by models, and features, it is important to consider these factors when investigating and comparing findings across different models.
Researchers have also investigated the impact of LLMs in Stack Overflow. Even before the release of LLMs, a decline in the number of answers to questions related to different programming languages was observed on Stack Overflow [24, 25, 72]. Recently, Rio-Chanona et al. [26] report an overall decrease in posting activity on Stack Overflow after the release of ChatGPT. However, whether such a decrease is uniform in different domains is unknown, as different topics involve different groups and their engagement. The decrease in users’ activity on Stack Overflow not only affects how developers receive support for their issues but also influences future versions of LLMs, given the need to train new models with updated data.
@@
Aiming to address the previous gaps, in this paper, we explore LLMs’ reliability, challenges, and impact on Stack Overflow, focusing on exploring different models and users’ activity regarding different topics and groups. To that end, we perform an empirical study analyzing questions and associated answers mined from Stack Overflow. First, we mine questions and related information from Stack Overflow (before and after ChatGPT’s release). Then, we analyze the overall and by topics impact of posting activity. Second, following previous studies conducted on Stack Overflow, we select a representative number of the previously selected questions to prompt LLMs, further comparing the generated answers with the original ones provided on Stack Overflow. To mitigate previous studies’ diversity issues regarding LLMs under analysis, we investigate two models: ChatGPT-3.5 and LLaMA-2-7b. Next, based on the answers provided, we further manually analyze the ones that LLMs present low similarity compared with human answers, reporting the most challenging topics they faced.
@@
We observe that ChatGPT-3.5 significantly outperforms LLaMA-2-7b regarding answers’ textual similarity. Although LLaMA-2-7b does not present significantly better results than ChatGPT, this model represents an alternative, free, and public option for the tech community. About the challenges faced by the LLMs, overall, they performed well in most domains, with some challenges regarding questions targeting Frameworks and Libraries, leading to different observed impacts on specific topics. Although ChatGPT presents a more neutral sentiment when generating answers, LLaMA presents a more positive sentiment. Finally, regarding the impact on Stack Overflow, our results confirm previous findings about a continuous significant decline in posting, answering, and commenting activity after ChatGPT’s release. Such a decline is not uniform initially, as we observed no statistical decline for some topics, like specific frameworks and libraries. However, after one year of ChatGPT’s release, we observe a constant and significant drop in posted content in Stack Overflow.
@@
  • 1.
    We report an empirical study evaluating the reliability of answers provided by ChatGPT and LLaMA in comparison to human-generated answers on Stack Overflow.
  • 2.
    We identify and further explore the domains where LLMs fall short and show that for some topics, like frameworks and libraries.
  • 3.
@@
  • 4.
    We reinforce previous findings about a decline in posting activity on Stack Overflow, though our results also show that such a decline is not uniform on different topics and groups.
  • 5.
    We provide a dataset of questions from Stack Overflow and associated scripts available through our online Appendix [27] and Zenodo as well. [^3] https://doi.org/10.5281/zenodo.15086541
@@ 2 Background
Stack Overflow: Stack Overflow (SO) is a platform where developers can ask questions and share their knowledge with others [28]. It is designed to provide accurate answers to specific programming problems, fostering a community of learning and sharing among developers. Over time, developers have adopted this platform to discuss different topics, while building general knowledge with the tech community.
Large Language Model: A large language model (LLM) is a machine learning model trained on extensive sets of types of data, like textual, images, audio, and structured data. Recently, LLMs trained on textual data brought the attention of users, scientists, and maintainers due to their support of generating human-like text [29]. With millions to billions of parameters, LLMs are used for various natural language processing tasks. Different LLMs have been proposed for tasks related to software engineering, like StarCoder [30], Copilot [31], and CodeBERT [32].
ChatGPT: ChatGPT is a language model developed by OpenAI. It is designed to generate human-like text based on the prompts it receives. Built on the GPT architecture, it can be used for a variety of applications, among which are code generation, summarization, translation, testing, and documentation [33, 34]. Currently, OpenAI offers different models, but in this study, we focus on ChatGPT-3.5. The motivation for this decision is driven by the recurrent adoption of this version by related work; furthermore, it is the default version that users have access to through the app option without further costs.
LLaMA: LLaMA, released by Meta AI, is a series of large language models that come with different sizes (e.g., 7, 13, and 70 billion parameters for LLaMA-2) [14]. Similar to ChatGPT, it spans various use cases, such as interacting with humans. In our study, we relied on the latest LLaMA version, i.e., LLaMA-2. This version has similarities with ChatGPT, like its architecture based on transformers and its generative capabilities, leading to generate coherent and contextually relevant text based on input prompts. Although this model has not been explored like ChatGPT, we aim to investigate how these models differ and complement each other. Specifically for this study, we consider LLaMA-2-7b as it represents the most accessible option for users, based on the required resources to load and use the model locally.
Although there are newer versions of the selected LLMs in this study, like GPT-4 and LLaMA 4, not evaluated here, we consider the selected versions based on their popularity and availability for the general users. Our study was conducted in two phases: first, we analyzed data before and after the release of ChatGPT, then, one year later, we collected data for the one-year milestone. When we conducted our study for the first phase in May 2023, GPT-3.5-turbo was the publicly available model. GPT-4 was released in March 2023 but was only accessible to (i) ChatGPT Plus subscribers or (ii) API users with a waitlist. This way, our analysis was necessarily based on ChatGPT-3.5, the model available to the general public at that time through its GUI without additional costs compared with ChatGPT-4. Regarding LLaMA, newer versions are known for handling more parameters and, consequently, dealing with more complex tasks. LLaMA API shares the same costs previously discussed, though their models can be locally loaded by users in their environments. However, it requires advanced resources due to hardware and software resource constraints, which might be challenging for some users.
@@ 3 Study Setup
Our experimental methodology investigates the reliability, challenges, and impact of LLMs on Stack Overflow. To this end, our study aims to answer the following research questions:
  • 1.
    RQ1. How does the reliability of answers generated by ChatGPT and LLaMA-2 compare to those provided by Stack Overflow users?
    Previous studies have investigated the adoption of LLMs to address Stack Overflow questions by evaluating the potential of ChatGPT for such a task [22, 23]. Knowing the diversity of LLMs available and their different configurations, we take a different route by evaluating the reliability of different LLMs while comparing them. Reliability in Software can be understood as the probability of having failure-free software execution for a timely and spacely specified period [35]. In this study, we investigate reliability based on the capability of LLMs to report answers matching the specifications reported in questions. Knowing that different answers can fix a single question in Stack Overflow, we focus on the answer previously accepted by the requester. To evaluate the reliability of LLM-generated answers, we use the accepted answer on Stack Overflow as a proxy, since it is typically considered the most reliable solution by the community, based on user feedback. We aim to investigate the capability of LLMs to report reliable and valuable content to users, considering the occurrence of wrong answers due to the LLM faults [36]. This way, we can evaluate how far or close LLMs are to the most accurate information that properly fixes the problem in that specific context. Following previous studies [37, 38], we randomly select a representative number of previously collected questions and prompt the models to answer them. Next, we compute the reliability by checking the textual and semantic similarity of the generated and original answers. Finally, knowing that answers avoiding negative comments are more likely to be accepted by Stack Overflow users [39], we aim to assess how close the generated answers fit with general users’ preferences. For that, we perform a sentiment analysis of the generated answers.
  • 2.
    RQ2. What specific domains pose challenges for ChatGPT and LLaMA-2 when generating reliable and accurate answers? When evaluating the adoption of LLMs to address Stack Overflow questions, previous studies’ findings report how well LLMs perform [22, 23]. However, there is a gap regarding how LLMs perform in different domains. Aiming to explore such a perspective, in this RQ, we investigate how to leverage LLMs’ adoption for practitioners. Based on the results of RQ1, for a subset of questions (low and high text similarity), we analyze and classify their associated domains, based on the categorization proposed by Barua et al. [28], reporting the domains that LLMs performed well and the challenging ones. Second, we explore the possible structural factors of the questions that correlate with high or low similarity.
  • 3.
    RQ3. What is the impact of ChatGPT ’s public release on Stack Overflow posting activity?
    • (a)
      RQ3.1. How are different domains from Stack Overflow impacted by the release of ChatGPT?
    Rio-Chanona et al. [26] and Burtch [40] investigate the impact of ChatGPT in Stack Overflow. Although they report a decline in user posting activities and Stack Overflow visits, it is unknown whether such a decline is uniform for different domains, like programming languages, frameworks, and libraries. Furthermore, they investigate the immediate impact without checking the evolution over time. Burtch et al. [40] even provide an analysis by individual topics but do not group related ones and analyze them together. Aiming to bring light to these concerns, we first mine questions and related information from Stack Overflow for five months before and after the release of ChatGPT (30th of November 2022), and after one year of the release (30th of November 2023). Second, we statistically assess the impact on the usage/engagement by comparing the number of posted questions, answers, comments, and related information. Third, to investigate the impact of ChatGPT on specific domains of Stack Overflow, we filter previously mined questions, which represent challenges for LLMs, and compare the frequency of posted questions referring to the associated tags. Based on our findings, we discuss possible factors that might influence the results and the different strategies for adopting/combining Stack Overflow and LLMs from now on.
  • 4.
@@
    RQ4. How has the release of ChatGPT impacted the activity patterns of Stack Overflow users? When investigating the impact of ChatGPT in Stack Overflow, Rio-Chanona et al. [26] mostly focus on the posting activity, which does not cover how such an impact affects the activity of users in Stack Overflow. Burtch et al. [40] and Xue et al. [41] even investigate the impact on users, focusing on the access traffic and the way users deal with Stack Overflow, respectively. In this RQ, we take a different route by investigating the impact of active users in Stack Overflow. First, we compare users’ activity level on Stack Overflow pre- and post-release of ChatGPT, considering posting questions, answers, and comments. Second, based on the different groups of users observed (reminiscent and new users), we further explore how users combine the adoption of ChatGPT and Stack Overflow. For that, we manually analyze questions not initially addressed by the LLM and later posted on the web forum. Based on our findings, we discuss the challenges Stack Overflow faces regarding their new users while understanding their new current needs.
@@
The remainder of this section describes the empirical setup of our study. We structure our study in three main steps, as presented in Figure 1: ❶ we mine SO questions using its API from July 2022 till May 2023, and Nov 2023 till May 2024; then, we filter the questions with accepted answers and sort them based on their score. ❷ we select a representative number of questions from this subgroup and use them to prompt ChatGPT and LLaMA-2; and ❸ we further collect metrics regarding question contents and associated information.